Martin has a total of 19 nickels and dimes worth $ 1.65. How many of each type of coin does he have?
Answer by ikleyn(52781)

.
     n +   d =  19            (1)
    5n + 10d = 165   cents    (2)



     n +   d =  19            (1')
     n +  2d =  33            (2')


Subtract equation (1') from (2').


           d = 33 - 19 = 14.


ANSWER.  14 dimes and 19-14 = 5 nickels.